two workers a and b together could finish the work in 8 days.they worked together for six days and a left the work.the remaining work was completed by b alone in 6 days.how many days would each take?

Answer by jorel1380(3719)

If we have two workers, a & b, finishing a job in 8 days, then:
1/a + 1/b=1/8
They both work six days, so:
6/a + 6/b=6/8=3/4
This leaves 1/4 of the work to be done, which b does in 6 days, so:
6/b=1/4
b=24
1/8-1/24=1/a
3/24-1/24=1/a
2a=24
a=12
a alone would take 12 days to finish the job, b would take 24 days working alone. ☺☺☺☺